# Fashion Image-to-Image Translation for Complementary Item Retrieval

### v2

@@ Abstract.
The increasing demand for online fashion retail has significantly boosted research in fashion compatibility modeling and item retrieval. Researchers focus on facilitating retrieval based on user queries, represented as textual descriptions or reference images, by modeling the compatibility among fashion items. Complementary item retrieval, especially top-bottom retrieval, presents challenges due to the need for precise compatibility modeling and traditional methods, predominantly based on Bayesian Personalized Ranking (BPR), often exhibit limited performance. More recently, researchers have focused on integrating generative models into compatibility modeling and complementary item retrieval systems to advance this field. These models produce item images that serve as additional inputs for the compatibility model. Interestingly, in the proposed approaches the quality of the generated images is often overlooked, although it could be crucial for the performance of the models. Moreover, although generative models can be effective, they usually require large amounts of data to achieve optimal performance, which can be problematic in contexts where such data is unavailable.
Aiming to overcome these limitations, we introduce the Generative Compatibility Model (GeCo), a two-stage approach that enhances fashion image retrieval through paired image-to-image translation. Initially, the Complementary Item Generation Model (CIGM), based on Conditional Generative Adversarial Networks, generates images of the target samples (e.g., bottoms) from seed items (e.g., tops), providing conditioning signals for retrieval. Subsequently, these samples are integrated into GeCo, refining compatibility modeling and retrieval accuracy. Our evaluations on three datasets demonstrate GeCo’s superior performance compared to state-of-the-art baselines. Our contributions include: (i) the GeCo model leveraging paired image-to-image translation within the Composed Image Retrieval framework, (ii) a comprehensive evaluations on benchmark datasets, and (iii) the release of a new Fashion Taobao dataset tailored for top-bottom retrieval, to foster further research.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Information retrievalccs: Information systems Multimedia and multimodal retrievalccs: Information systems Recommender systems
@@ 1. Introduction
Retrieving fashion items that go together well necessitates a robust approach capable of evaluating the compatibility of two items. Classical approaches often rely on the Bayesian Personalized Ranking (BPR) framework (Song et al., 2017), leveraging visual and textual information. BPR allows the optimization of the ranking of items by maximizing the difference in predicted scores between a positive and a negative item pair, tailored specifically for implicit feedback scenarios. Even though the integration of generative models in top-bottom retrieval has proven to be effective (Liu et al., 2020a, b; Lin et al., 2019), these models are primarily optimized for compatibility modeling and often produce low-quality images of small sizes. Furthermore, generative architecture-based models require significant amounts of data to ensure that they accurately capture the underlying distribution and compatibility relationships. This makes it difficult to effectively use such models when working with small datasets.
@@
In the second stage, GeCo incorporates these generated samples to perform compatibility modeling and complementary item retrieval. Specifically, we model these tasks within the Composed Image Retrieval (CoIR) framework, which is dedicated to retrieve images using queries composed of multiple elements (Baldrati et al., 2024; Feng et al., 2023; Liu et al., 2021; Feng et al., 2024). While previous works model the query by combining an input image and a textual description of the desired outcomes, our approach formulates the query by considering both the top and the template images. This results in a more precise modeling of item compatibility and improved retrieval performance. The structure of our model is depicted in Figure 1. By placing greater importance on the generated images compared to previous works in the fashion domain (Liu et al., 2020a, b; Lin et al., 2019), this formulation motivates us to rely solely on visual information for the CIGM and GeCo models, distinguishing our approach from prior studies.
Moreover, our proposed two-stage architecture is designed to be efficient and requires less training data compared to existing approaches, due to the absence of the end-to-end training of the two models. To the best of our knowledge, while prior works in CoIR have predominantly focused on creating joint queries combining images with textual descriptions (Baldrati et al., 2023, 2022), our work is the first to model CoIR using top images and templates as queries, performing compatibility modeling and complementary item retrieval.
@@
In the remainder of this paper, we first review related work in fashion image retrieval and generative models. Then, we provide the technical background on generative models, needed for understanding the details of the proposed approach, including the architecture and training methodology. In Section 5, we present the experimental setup and results, highlighting the performance improvements achieved by our method. Finally, we conclude the paper and discuss potential future research directions.
@@ 2.2. Generative Models for Top-Bottom Retrieval
Notable contributions in the domain of complementary item retrieval include the work by Kang et al. (2017), who proposed DVBPR, a method that incorporates a Siamese Network and a Generative Adversarial Network (GAN) for compatibility modeling and generation. This method is augmented with a preference maximization module that also includes personalization. Other approaches, such as FARM (Lin et al., 2019) and MGCM (Liu et al., 2020a), generate novel bottoms using advanced architectures. FARM integrates a variational transformer to produce compatible bottoms given a top, leveraging learned latent spaces for compatibility modeling. In contrast, MGCM introduces a compatibility schema by harnessing a GAN to generate bottom templates and evaluates the compatibility of top-bottom pairs. Specifically, MGCM considers both item-item and item-template relationships: item-item compatibility refers to the compatibility between the top and candidate bottom features, while item-template compatibility denotes the similarity between the candidate and the template. The inclusion of templates in the compatibility modeling enhances both the compatibility modeling and retrieval task performance. Following a similar approach, Liu et al. (2020b) propose to assess item-item and item-template compatibilities by employing a Cycle-GAN (Zhu et al., 2017) generator for the compatibility module.
@@ 3. Technical Background: Generative Models
In this section, we provide the essential technical background to understand the design choices behind our approach. We begin by defining generative models, followed by a detailed description of three pivotal ones: Variational Autoencoders, Generative Adversarial Networks, and Stable Diffusion. We discuss the characteristics and limitations of each of these approaches and clarify our selection of GAN.
@@ 3.1. Variational Autoencoders
The first remarkable work is represented by the introduction of Variational Autoencoders (VAEs) (Kingma and Welling, 2014). VAEs consist of an encoder, denoted as $q_{\phi}(z|x)$, which maps input data $x$ to a latent space representation $z$. This encoder approximates the posterior distribution $p(z|x)$ using a variational distribution $q_{\phi}(z|x)$, parameterized by $\phi$. The training objective involves maximizing the evidence lower bound (ELBO). The decoder, represented as $p_{\theta}(x|z)$, generates samples $x$ from the latent space $z$, effectively mapping $z$ back to the data space. Although VAEs have demonstrated its effectiveness in various image generation tasks, they often produce blurry samples due to the minimization of the Kullback-Leibler (KL) divergence between the sampled latent representations and the model’s prior distribution (Goodfellow, 2017), as illustrated in Figure 2(a).
@@ 3.2. Generative Adversarial Networks
(1) $\min_{G}\max_{D}V(D,G)=\mathbb{E}_{\mathbf{x}\sim p_{data}(\mathbf{x})}[\log D(\mathbf{x})]+\mathbb{E}_{\mathbf{z}\sim p_{\mathbf{z}}(\mathbf{z})}[\log(1-D(G(\mathbf{z})))]$
@@
(2) $\min_{G}\max_{D}V(D,G)=\mathbb{E}_{\mathbf{x},\mathbf{y}\sim p_{data}(\mathbf{x},\mathbf{y})}[\log D(\mathbf{x},\mathbf{y})]+\mathbb{E}_{\mathbf{x}\sim p_{data}(\mathbf{x}),\mathbf{z}\sim p_{\mathbf{z}}(\mathbf{z})}[\log(1-D(\mathbf{x},G(\mathbf{x},\mathbf{z})))],$
@@
While GAN models are capable of effectively tackling and solving complex tasks via adversarial learning, such objective is a double-edged sword: while it drives the generator to improve continually, it can also lead to instability. If the discriminator becomes too strong, it can easily distinguish between real and fake samples, providing little information to the generator; conversely, if the generator becomes too strong, the discriminator may fail to provide meaningful feedback (Miyato et al., 2018). This imbalance can cause the generator to focus on a narrow subset of the data distribution that consistently fools the discriminator, leading to mode collapse (Thanh-Tung and Tran, 2020; Mescheder et al., 2018; Arjovsky and Bottou, 2017). This occurs when the generator learns few modes of the original data distribution that actually fool the discriminator, thus leading the training to divergence. The original GAN loss formulation, which is based on Jensen-Shannon (JS) divergence, works well when the distributions of real and generated data have little overlap. Alternative loss functions, such as those based on the Wasserstein distance (Arjovsky et al., 2017), have been proposed to address this issue by providing smoother gradients, which can help mitigate mode collapse to some extent. This loss leads to more stable training dynamics, but, employing the Wasserstein distance necessitates that the discriminator be Lipschitz-1 continuous. To enforce this condition, weight clipping was initially used. Subsequently, it was proven that gradient penalty (Gulrajani et al., 2017) or spectral normalization (Miyato et al., 2018) could also be effectively employed to maintain the Lipschitz constraint.
@@ 3.3. Stable Diffusion
We note that due to certain intrinsic limitations of adversarial training, researchers focused their attention on other generative models. In particular, Diffusion Models (Rombach et al., 2022; Ho et al., 2020) gained importance due to the highly realistic samples that such models can generate, which according to Dhariwal and Nichol (2021) is superior compared to GANs while retaining stable training. Diffusion models work with the forward and the backward processes. The forward process gradually corrupts the input pattern by injecting noise according to a specific Stochastic Differential Equation (SDE) process, fixed at different timesteps. In contrast, the backward process attempts to estimate the noise injected at each timestep using a complex neural network. During this backward process, the estimated noise is progressively removed from the corrupted pattern to reconstruct the original input, thereby approximating another SDE. Recent works have proposed estimating the SDE via a UNet architecture (Ronneberger et al., 2015). Additionally, various samplers can be employed for generating samples from the SDE learnt by the Diffusion Model (Karras et al., 2022; Song et al., 2021a; Ho et al., 2020). Depending on the chosen fixed forward process SDE and the number of timesteps, several strategies have been developed to enhance performance (Song et al., 2021a, b; Liu et al., 2022).
Despite their high-quality generations, Diffusion Models have yet to see widespread application within the fields of information retrieval and recommender systems. This contrasts VAEs and GANs, which have been extensively explored and integrated into these domains. The primary barrier to adopting Diffusion Models is the significant computational cost associated with their architectures (Rombach et al., 2022; Berthelot et al., 2024). These models require powerful computational resources for training and inference due to their iterative nature, which involves multiple steps of noise injection and removal through complex neural networks, thus increasing the computational resources required compared to other generative models. As a result, despite their superior ability to generate realistic samples, the practical deployment of Diffusion Models in real-time applications remains challenging. The high computational overhead limits their integration into systems where efficiency and rapid response times are essential.
@@ 4.1. Problem Formulation
We recall that we propose a methodology that addresses compatibility modeling and complementary item retrieval tasks using a generative model. To achieve this, we need to model the probability distributions over the sets of tops $\mathcal{T}$ and bottoms $\mathcal{B}$. Our approach first involves learning a mapping between the probability distributions of tops $P_{data}(\mathcal{T})$ and bottoms $P_{data}(\mathcal{B})$ using the CIGM model. Then we use this mapping to generate meaningful templates given a top $t\in\mathcal{T}$. Ideally, the CIGM model should learn to generate samples $b\in\mathcal{B}$. However, in practice, CIGM learns a probability distribution $P_{\textsc{CIGM}}(\mathcal{B})$ that approximates the probability distribution over the bottom set, such that $P_{\textsc{CIGM}}(\mathcal{B})\approx P_{data}(\mathcal{B})$. Using the generated template, we employ the GeCo model to assess the top-bottom compatibility.
Formally, given a dataset, we define $\mathcal{P}$ as the set of all compatible pairs $(t_{i},b_{j})$, with $t_{i}\in\mathcal{T}$ and $b_{j}\in\mathcal{B}$, and $\mathcal{N}$ as the set of all incompatible pairs $(t_{i},b_{k})$, with $b_{k}\in\mathcal{B}$. It is noteworthy that we consider only the pairs reported in the training data as compatible, treating all other combinations as negative examples. We first train the CIGM model to generate the bottom $\hat{b}_{i}$ corresponding to the input top $t_{i}$, following the Equation 3. Once the CIGM model is trained, we train the GeCo model to learn hidden compatibility relations. This is accomplished by considering the projection onto a fixed latent space of the top $t_{i}$ and the template $\hat{b}_{i}$, which are referred to as the query. The model learns to align the query latent space with the bottom one, thereby encouraging compatible elements to exhibit more remarkable similarity, as detailed in the Equation 4. We first explain the model used to create templates and then discuss the GeCo model.
@@ 4.2. Complementary Item Generation Model (CIGM)
As foreshadowed in the previous section, in order to overcome the mode collapse phenomenon (Thanh-Tung and Tran, 2020; Mescheder et al., 2018) and to produce more realistic templates, we modified the generator architecture to incorporate noise at the bottleneck level. Specifically, the latent representation is concatenated with a noise vector $\mathbf{z}\sim p(\mathbf{z})$, where $p(\mathbf{z})$ is a Gaussian distribution. This noise injection aims to add stochasticity to the generation process, thereby increasing the diversity of the generated samples, while maintaining the simplicity inherent in a paired image-to-image translation model. Additionally, the skip connections are retained to leverage the benefits of the UNet architecture for better feature propagation during training. Figure 4 illustrates the adapted architecture.
@@
(6) $\min_{G}\max_{D}\mathcal{L}_{\text{GAN}}=\mathbb{E}_{\mathbf{t_{i}}\sim\mathcal{T},\mathbf{b_{j}}\sim\mathcal{B}}[\log D(\mathbf{t_{i},b_{j}})]+\mathbb{E}_{t_{i}\sim\mathcal{T},\mathbf{z}\sim p_{\mathbf{z}}(\mathbf{z})}[\log(1-D(t_{i},G(\mathbf{t_{i},z})))]$
(7) $\mathcal{L}_{G}=-\mathbb{E}_{t_{i}\sim\mathcal{T},\mathbf{z}\sim p_{\mathbf{z}}(\mathbf{z})}[\log D(t_{i},G(t_{i},\mathbf{z}))]+\lambda\|G(t_{i},\mathbf{z})-b_{j}\|_{1}$
(8) $\mathcal{L}_{D}=-\mathbb{E}_{\mathbf{t_{i}}\sim\mathcal{T},\mathbf{b_{j}}\sim\mathcal{B}}[\log D(\mathbf{t_{i},b_{j}})]-\mathbb{E}_{\mathbf{t_{i}}\sim\mathcal{T},\mathbf{z}\sim p_{\mathbf{z}}(\mathbf{z})}[\log(1-D(t_{i},G(\mathbf{t_{i},z})))]$
@@
Given the min-max game expressed in Equation 6, the generator aims to minimize the objective in Equation 7, where $\|G(t_{i},\mathbf{z})-b_{j}\|_{1}$ is a regularization term that guides the generator, weighted by the hyperparameter $\lambda$. Conversely, the discriminator aims to minimize the objective in Equation 8. The goal is to reach a Nash equilibrium where the generator successfully fools the discriminator. Once this equilibrium is achieved, only the generator is employed in the proposed GeCo model. The discriminator’s role is solely to guide the generator in learning the underlying distribution. Figure 5 displays some samples using the modified UNet generator with added noise. As depicted in Figure 5(c), the generated bottom images are not only meaningful but also exhibit diversity compared to the actual ground-truth shown in Figure 5(b).

### v3

@@ Abstract.
The increasing demand for online fashion retail has significantly boosted research in fashion compatibility modeling and item retrieval. Researchers focus on facilitating retrieval based on user queries, represented as textual descriptions or reference images, by modeling the compatibility among fashion items. Complementary item retrieval, especially top-bottom retrieval, presents challenges due to the need for precise compatibility modeling and traditional methods, predominantly based on Bayesian Personalized Ranking, often exhibit limited performance. More recently, researchers have focused on integrating generative models into compatibility modeling and complementary item retrieval systems to advance this field. These models produce item images that serve as additional inputs for the compatibility model. Interestingly, in the proposed approaches the quality of the generated images is often overlooked, although it could be crucial for the performance of the models. Moreover, although generative models can be effective, they usually require large amounts of data to achieve optimal performance, which can be problematic in contexts where such amount can be limited.
Aiming to overcome these limitations, we introduce the Generative Compatibility Model (GeCo), a two-stage approach that enhances fashion image retrieval through paired image-to-image translation. Initially, the first-stage model, called Complementary Item Generation Model, which is based on Conditional Generative Adversarial Networks, generates images of the target samples (e.g., bottoms) from seed items (e.g., tops), providing conditioning signals for retrieval. Subsequently, these samples are integrated into GeCo, the second stage, refining compatibility modeling and retrieval accuracy. Overall, our evaluations on three datasets demonstrate GeCo’s superior performance compared to state-of-the-art baselines. Our contributions include: (i) the GeCo model leveraging paired image-to-image translation within the Composed Image Retrieval framework, (ii) a comprehensive evaluations on benchmark datasets, and (iii) the release of a new Fashion Taobao dataset tailored for top-bottom retrieval, to foster further research. Our code is available at: https://github.com/sisinflab/GeCo.
@@
conference: Pre-print; under review; September 2024copyright: noneccs: Information systems Information retrievalccs: Information systems Multimedia and multimodal retrievalccs: Information systems Recommender systems
@@ 1. Introduction
Retrieving fashion items that go together well necessitates a robust approach capable of evaluating the compatibility of two items. Classical approaches often rely on the Bayesian Personalized Ranking (BPR) framework (Song et al., 2017), leveraging visual and textual information. BPR allows the optimization of the ranking of items by maximizing the difference in predicted scores between a positive and a negative item pair, tailored specifically for implicit feedback scenarios. Even though the integration of generative models in top-bottom retrieval has proven to be effective (Liu et al., 2020a, b; Lin et al., 2019), these models are primarily optimized for compatibility modeling and often produce low-quality images of small sizes. Furthermore, generative architecture-based models may require significant amounts of data to ensure that they accurately capture the underlying distribution and compatibility relationships. This makes it difficult to effectively use such models when working with small datasets.
@@
In the second stage, GeCo incorporates these generated samples to perform compatibility modeling and complementary item retrieval. Specifically, we model these tasks within the Composed Image Retrieval (CoIR) framework, which is designed to retrieve images using queries composed of multiple elements (Baldrati et al., 2024; Feng et al., 2023; Liu et al., 2021; Feng et al., 2024). While previous works model the query by combining an input image and a textual description of the desired outcomes, our approach formulates the query by considering both the top and the template images. This results in a more precise modeling of item compatibility and improved retrieval performance. The structure of our model is depicted in Figure 1. By placing greater importance on the generated images compared to previous works in the fashion domain (Liu et al., 2020a, b; Lin et al., 2019), this formulation motivates us to rely solely on visual information for the CIGM and GeCo models, distinguishing our approach from prior studies. Moreover, our proposed two-stage architecture is designed to be efficient and requires less training data compared to existing approaches, because it does not require an end-to-end-training. To the best of our knowledge, while prior works in CoIR have predominantly focused on creating joint queries combining images with textual descriptions (Baldrati et al., 2023, 2022), our work is the first to model CoIR using top images and templates as queries, performing compatibility modeling and complementary item retrieval.
@@
In the remainder of this paper, we review related work in fashion image retrieval and generative models. Next, we provide the technical background on generative models, focusing on the architecture and training methodology relevant to understanding the proposed approach. After that, we present the experimental setup and results, emphasizing the performance improvements achieved by our method. Finally, we discuss the paper’s findings and limitations and sketch future research directions.
@@ 2.2. Generative Models for Top-Bottom Retrieval
Notable contributions in the domain of complementary item retrieval include the work by Kang et al. (2017), who proposed DVBPR, a method that incorporates a Siamese Network and a Generative Adversarial Network (GAN) for compatibility modeling and generation. This method is augmented with a preference maximization module that also includes personalization. Other approaches, such as FARM (Lin et al., 2019) and MGCM (Liu et al., 2020a), generate novel bottoms using advanced architectures. FARM integrates a variational transformer to produce compatible bottoms given a top, leveraging learned latent spaces for compatibility modeling. In contrast, MGCM introduces a compatibility schema by harnessing a GAN to generate bottom templates, including them in the compatibility assessment. Specifically, MGCM considers both item-item and item-template relationships: item-item compatibility refers to the compatibility between the top and candidate bottom features, while item-template compatibility denotes the similarity between the candidate and the template. The inclusion of templates in the compatibility modeling enhances both the compatibility modeling and retrieval task performance. Following a similar approach, Liu et al. (2020b) propose to assess item-item and item-template compatibilities by employing a Cycle-GAN (Zhu et al., 2017) generator for the compatibility module.
@@ 3. Technical Background: Generative Models
In this section, we provide the essential technical background to understand the design choices behind our approach. We begin by defining generative models, followed by a detailed description of three pivotal ones: Variational Autoencoders, Generative Adversarial Networks, and Stable Diffusion. We discuss the characteristics and limitations of these approaches, which led us to the choice of GANs for our method.
@@ 3.1. Variational Autoencoders
The first remarkable work is represented by the introduction of Variational Autoencoders (VAEs) (Kingma and Welling, 2014). VAEs consist of an encoder, denoted as $q_{\phi}(z|x)$, which maps input data $x$ to a latent space representation $z$. This encoder approximates the posterior distribution $p(z|x)$ using a variational distribution $q_{\phi}(z|x)$, parameterized by $\phi$. The training objective involves maximizing the evidence lower bound (ELBO). The decoder, represented as $p_{\theta}(x|z)$, generates samples $x$ from the latent space $z$, effectively mapping $z$ back to the data space. Although VAEs have demonstrated their effectiveness in various image generation tasks, they often produce blurry samples due to the minimization of the Kullback-Leibler (KL) divergence between the sampled latent representations and the model’s prior distribution (Goodfellow, 2017), as illustrated in Figure 2(a).
@@ 3.2. Generative Adversarial Networks
(1) $\min_{G}\max_{D}V(D,G)=\mathbb{E}_{x\sim p_{data}(x)}[\log D(x)]+\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))]$
@@
(2) $\min_{G}\max_{D}V(D,G)=\mathbb{E}_{x,y\sim p_{data}(x,y)}[\log D(x,y)]+\mathbb{E}_{x\sim p_{data}(x),z\sim p_{z}(z)}[\log(1-D(x,G(x,z)))],$
@@
While GAN models are capable of effectively tackling and solving complex tasks via adversarial learning, such objective is a double-edged sword: while it drives the generator to improve continually, it can also lead to instability. If the discriminator becomes too strong, it can easily distinguish between real and fake samples, providing little information to the generator; conversely, if the generator becomes too strong, the discriminator may fail to provide meaningful feedback (Miyato et al., 2018). This imbalance can cause the generator to focus on a narrow subset of the data distribution that consistently fools the discriminator, leading to mode collapse (Thanh-Tung and Tran, 2020; Mescheder et al., 2018; Arjovsky and Bottou, 2017). This occurs when the generator learns few modes of the original data distribution that actually fool the discriminator, thus leading the training to divergence. The original GAN loss formulation, which is based on Jensen-Shannon (JS) divergence, works well when the distributions of real and generated data have little overlap. Alternative loss functions, such as those based on the Wasserstein distance (Arjovsky et al., 2017), have been proposed to address this issue by providing smoother gradients, which can help mitigate mode collapse to some extent. This loss leads to more stable training dynamics, but, employing the Wasserstein distance necessitates that the discriminator be Lipschitz-1 continuous. To enforce this condition, weight clipping was initially used. Subsequently, it was shown that gradient penalty (Gulrajani et al., 2017) or spectral normalization (Miyato et al., 2018) could also be effectively employed to maintain the Lipschitz constraint.
@@ 3.3. Stable Diffusion
Given the inherent limitations of adversarial training discussed earlier, researchers shifted their focus to other generative models. In particular, Diffusion Models (Rombach et al., 2022; Ho et al., 2020) gained importance due to the highly realistic samples that such models can generate, which outperforms GANs (Dhariwal and Nichol, 2021) while retaining stable training. Diffusion models work with the forward and the backward processes. The forward process gradually corrupts the input pattern by injecting noise according to a specific Stochastic Differential Equation (SDE) process, fixed at different timesteps. In contrast, the backward process attempts to estimate the noise injected at each timestep using a complex neural network. During this backward process, the estimated noise is progressively removed from the corrupted pattern to reconstruct the original input, thereby approximating another SDE. Recent works have proposed estimating the SDE via a UNet architecture (Ronneberger et al., 2015). Additionally, various samplers can be employed for generating samples from the SDE learnt by the Diffusion Model (Karras et al., 2022; Song et al., 2021a; Ho et al., 2020). Depending on the chosen fixed forward process SDE and the number of timesteps, several strategies have been developed to enhance performance (Song et al., 2021a, b; Liu et al., 2022).
Despite their high-quality generations, Diffusion Models have yet to see widespread application within the fields of information retrieval and recommender systems. This contrasts VAEs and GANs, which have been extensively explored and integrated into these domains. The primary barrier to adopting Diffusion Models is the significant computational cost associated with their architectures (Rombach et al., 2022; Berthelot et al., 2024). These models require powerful computational resources for training and inference due to their iterative nature, which involves multiple steps of noise injection and removal through complex neural networks, thus increasing the computational resources required compared to other generative models. As a result, despite their superior ability to generate realistic samples, the practical deployment of Diffusion Models in real-time applications remains challenging. The high computational overhead limits their integration into systems where efficiency and rapid response times are essential, such as CoIR systems, where rapid responsiveness is crucial (Liu et al., 2021).
@@ 4.1. Problem Formulation
We recall that we present a methodology that addresses compatibility modeling and complementary item retrieval tasks using a generative model. To achieve this, we need to model the probability distributions over the sets of tops $\mathcal{T}$ and bottoms $\mathcal{B}$. Our approach first involves learning a mapping between the probability distributions of tops $P_{data}(\mathcal{T})$ and bottoms $P_{data}(\mathcal{B})$ using the CIGM model. Then we use this mapping to generate meaningful templates given a top $t\in\mathcal{T}$. Ideally, the CIGM model should learn to generate samples $b\in\mathcal{B}$. However, in practice, CIGM learns a probability distribution $P_{\textsc{CIGM}}(\mathcal{B})$ that approximates the probability distribution over the bottom set, such that $P_{\textsc{CIGM}}(\mathcal{B})\approx P_{data}(\mathcal{B})$. Using the generated template, we employ the GeCo model to assess the top-bottom compatibility.
Formally, given a dataset, we define $\mathcal{P}$ as the set of all compatible pairs $(t_{i},b_{j})$, with $t_{i}\in\mathcal{T}$ and $b_{j}\in\mathcal{B}$, and $\mathcal{N}$ as the set of all incompatible pairs $(t_{i},b_{k})$, with $b_{k}\in\mathcal{B}$. It is noteworthy that we consider only the pairs reported in the training data as compatible, treating all other combinations as negative examples. We first train the CIGM model to generate the bottom $\hat{b}_{i}$ corresponding to the input top $t_{i}$, following Equation 3. Once the CIGM model is trained, we train the GeCo model to learn hidden compatibility relations. This is accomplished by considering the projection onto a fixed latent space of the top $t_{i}$ and the template $\hat{b}_{i}$, which are referred to as the query. The model learns to align the query latent space with the bottom one, thereby encouraging compatible elements to be more similar, as detailed in the Equation 4. We first explain the model used to create templates and then discuss the GeCo model.
@@ 4.2. Complementary Item Generation Model (CIGM)
As foreshadowed in the previous section, in order to overcome the mode collapse phenomenon (Thanh-Tung and Tran, 2020; Mescheder et al., 2018) and to produce more realistic templates, we modified the generator architecture to incorporate noise at the bottleneck level. Specifically, the latent representation is concatenated with a noise vector $\mathbf{z}\sim p(\mathbf{z})$, where $p(\mathbf{z})$ is a Gaussian distribution. This noise injection aims to add stochasticity to the generation process, thereby increasing the diversity of the generated samples, while maintaining the simplicity inherent to a paired image-to-image translation model. Additionally, the skip connections are retained to leverage the benefits of the UNet architecture for better feature propagation during training. Figure 4 illustrates the adapted architecture.
@@
(6) $\min_{G}\max_{D}\mathcal{L}_{\text{GAN}}=\mathbb{E}_{t_{i}\sim\mathcal{T},b_{j}\sim\mathcal{B}}[\log D(t_{i},b_{j})]+\mathbb{E}_{t_{i}\sim\mathcal{T},z\sim p_{z}(z)}[\log(1-D(t_{i},G(t_{i},z)))]$
(7) $\mathcal{L}_{G}=-\mathbb{E}_{t_{i}\sim\mathcal{T},z\sim p_{z}(z)}[\log D(t_{i},G(t_{i},z))]+\lambda\|G(t_{i},z)-b_{j}\|_{1}$
(8) $\mathcal{L}_{D}=-\mathbb{E}_{t_{i}\sim\mathcal{T},b_{j}\sim\mathcal{B}}[\log D(t_{i},b_{j})]-\mathbb{E}_{t_{i}\sim\mathcal{T},z\sim p_{z}(z)}[\log(1-D(t_{i},G(t_{i},z)))]$
@@
Given the min-max game expressed in Equation 6, the generator aims to minimize the objective in Equation 7, where $\|G(t_{i},\mathbf{z})-b_{j}\|_{1}$ is a regularization term that guides the generator, weighted by the hyperparameter $\lambda$. Conversely, the discriminator aims to minimize the objective in Equation 8. The goal is to reach a Nash equilibrium where the generator successfully fools the discriminator. Once this equilibrium is achieved, only the generator is employed in the proposed GeCo model. The discriminator’s role is solely to guide the generator in learning the underlying distribution. Figure 5 displays some samples using the modified UNet generator with added noise. As depicted in Figure 5(c), the generated bottom images are not only realistic but also exhibit diversity compared to the actual ground-truth shown in Figure 5(b).